# A GNN-based Spectral Filtering Mechanism for Imbalance Classification in Network Digital Twin

### v1

@@ Abstract
Graph Neural Networks are gaining attention in Fifth-Generation (5G) core network digital twins, which are data-driven complex systems with numerous components. Analyzing these data can be challenging due to rare failure types, leading to imbalanced classification in multiclass settings. Digital twins of 5G networks increasingly employ graph classification as the main method for identifying failure types. However, the skewed distribution of failure occurrences is a major class imbalance issue that prevents effective graph data mining. Previous studies have not sufficiently tackled this complex problem. In this paper, we propose Class-Fourier Graph Neural Network (CF-GNN) introduces a class-oriented spectral filtering mechanism that ensures precise classification by estimating a unique spectral filter for each class. We employ eigenvalue and eigenvector spectral filtering to capture and adapt to variations in the minority classes, ensuring accurate class-specific feature discrimination, and adept at graph representation learning for complex local structures among neighbors in an end-to-end setting. Extensive experiments have demonstrated that the proposed CF-GNN could help with both the creation of new techniques for enhancing classifiers and the investigation of the characteristics of the multi-class imbalanced data in network digital twin system.
@@
Graph neural network , Multi-class imbalanced , Eigenvalue , Graph representation learning , Network Digital Twin
journal: Internet of Things
@@
We proposed a novel CF-GNN to address class imbalance tasks in multiclass 5G network digital twins.
CF-GNN employs the concepts of eigenvalues and eigenvectors to identify minority classes in a multiclass failure state.
Incorporating spectral filtering influenced by the global graph structure and local neighborhood information during the filtering process.
The proposed model showcased the robust performance of the network digital twin-unbalanced datasets.
@@ 1 Introduction
The emergence of fifth-generation (5G) networks, and robust core network systems are paramount for the evolution to Industry 5.0. According to recent studies, 5G connections are projected to increase over 100-fold, from approximately 13 million in 2018 to 2.6 billion by 2025 [1]. This growth is expected to continue with sixth-generation (6G) networks and the advent of the Industry 5.0 era, characterized by the vision of the Internet of Everything. In this context, network failure classification for 5G and beyond is essential to optime the network performance. As the backbone of communication infrastructure, reliable and resilient networks are essential for seamless operation in diverse applications in environments ranging from daily life to industrial settings. The core network management center can identify and classify network failures by applying deep learning algorithms and advanced analytics, enabling preemptive measures to ensure uninterrupted service delivery [2].
The digital twin (DT) is a powerful technology that has emerged as a promising solution for connecting physical spaces with digital systems [3, 4]. This technology uses historical data and real-time operational information to create digital replicas of physical entities, such as devices, machines, and objects. Research on DTs [5] and [6] has garnered significant attention, with numerous researchers exploring its applications in such fields as aviation, 6G networks, intelligent manufacturing, and data generation [7, 8]. Additionally, DT has been recognized as one of the 12 representative use cases for future networks by the International Telecommunication Union Network 2030 focus group [9]. However, setting up a replica of a network event to mimic every potential pattern in network DTs is challenging.
Numerous tasks related to graphs classification have been extensively studied, including node classification [10], graph classification [11], and link predictions [12]. Moreover, a variety of advanced GNN methodologies, such as graph pooling [13], graph convolution [14], and graph attention mechanisms [15], have been suggested to enhance the performance of GNNs. Despite these advancements, there remains a scarcity of graph models for communication network domains [16], which is crucial for understanding GNNs. Several methods, such as RouteNet [17] and RouteNet-Fermi [18], have been introduced to model the NDT for network performance. The benefit of data-driven models [19] is their ability to accurately capture a wide range of complex network nodes with unprecedented precision, as they are trained on real-world data [20]. Nevertheless, none of these recent studies have focused on data-driven network digital twin approach to tackle the imbalance node classification tasks in the NDT domain.
Class-Imbalanced across many application fields, including the Digital Twin [21], Fraud Detection [22], for computing [23], and recently the synthetic oversampling [24]. GNN models that are robust to class imbalance must be developed to avoid biases towards majority classes while maintaining the ability to generalize over minority classes. Traditional techniques for addressing the class imbalance, including oversampling [25], and sampling the classifiers [26], have limitations when applied to graph-structured data because they do not account for the inherent graph structure. Thus, many scenarios have been developed to tackle class imbalance within semi-supervised node classification. Node classification with synthetic oversampling methods, GraphSMOTE [27, 28], and GATE-GNN [29] generates new nodes and employ a variety of methods to connect them to the existing graph. Furthermore, these methods often fail to fully utilize the untapped potential of the abundant labeled and unlabeled nodes in the graph, which limits their effectiveness in real-world applications.
In this study, we present a novel Class-Fourier Graph Neural Network filtering (CF-GNN) framework designed to address the challenges posed by Imbalanced node classification tasks. Our proposed method leverage node-specific spectral filtering to ensure that minority-class nodes receive attention without being overtaken by majority-class nodes. By enabling the filter to focus on particular nodes. This helps to emphasize overshadowed classes, which is particular beneficial for imbalanced datasets. the localized filtering concentrates on the distinct features of nodes. This ensures that class-specific changes can be applied to nodes that belong to minority classes, enhancing their representation in training and prediction. Additionally, the filtering process, being aware of class-specific characteristics in the spectral domain, reduces the likelihood of minority-class nodes being overwhelmed by dominant classes. During training, the model employs class-weighted loss functions, which penalize the misclassification of minority-class samples more heavily. This forces the network to learn better representations of underrepresented classes. The spectral filtering procedure of the CF-GNN explicitly integrates a multiclass structure, transforming features in a class-aware way by the use of spectral coefficients and eigenvalue-based distinctions between classes. Our experimental evaluations on two 5G NDT datasets demonstrate how well the proposed CF-GNN framework compared to state-of-the-art GNN techniques designed to mitigate class imbalance problems.
@@
    The proposed CF-GNN employs a spectral attention technique in the Fourier domain to extract complex structural components that conventional spatial-based GNN may overlook.
@@
    CF-GNN utilizes class-specific spectral filtering to enhance classification accuracy by highlighting significant feature components associated with certain classes. By iteratively improving the eigenvector estimates, this procedure ensure accurate spectral analysis of the graph structure.
@@
    Introduces the theoretical Twin-GFT for multidimensional spectral representation, enabling scalable failure detection and improving prediction robustness through ensemble classification.
@@ 1.1 Related work
Graph learning-driven failure predictions have been gaining attention due to their effectiveness in tackling the massive volumes of data generated by 5G networks and their ability to understand the complex structural characteristics of these networks. This section explores three areas of research: one focusing on learning-based 5G failure prediction, imbalanced learning techniques, and Graph Fourier transform (GFT) to capture the complex structure of NDTs.
@@ 1.1.1 Learning-based failure prediction
The primary advantage of applying deep graph learning to network modeling [30] lies in its data-driven approach, which enable it to capture the complex nature of the real-world networks. Most existing research in [31] and [32] uses fully connected conventional neural networks. However, the primary constraint of these methods is their non-generalization to alternative network topologies and configurations, such as routing. In this context, more recent studies have proposed advanced neural network models, such as GNNs [33, 34], convolutional neural networks [32], and variational autoencoders [35] for robustness in imbalanced datasets common to cybersecurity domains. Despite their advancements, these models have different objectives and overlook critical aspects of real-world networks from the twin perspective. The concept of broadly applying neural networks to graphs has gained significant attention. For example, convolutional networks have been extended to graphs in the spectral domain [36], where filters are applied to frequency modes using GFT. The eigenvector matrix of a graph Laplacian must be multiplied to achieve this transformation. One study [37] parameterized the spectrum filters as Chebyshev polynomials of eigenvalues, resulting in efficient and localized filters and reducing the computational burden. However, a drawback of these spectral formulations are disadvantageous is that they are limited to graphs with a single structure since they depend on the fixed spectrum of the graph Laplacian. In contrast, spatial formulations are not constrained by a specific graph topology. Generalizing neural networks to graphs is studied in two ways: a) given a single graph structure, or labels of individual nodes [38, 36, 39, 40, 37, 41] b) given a set of graphs with different structure and sizes, the goal is learn predictions of the class labels of the graphs are learned [42, 43, 44]. Inspired by the ITU Challenge ”ITU-ML5G-PS-008: Network Failure Classification Model Using the Digital Twin Network” [45], our research aims to leverage GNNs, enhanced with the CF-GNN, to analyze interconnected systems and networks in an end-to-end manner. By integrating class-oriented spectral filtering, the proposed CF-GNN framework enables precise classification of network failure types and accurate identification of failure points. The approach improves accuracy and adaptability compared to conventional deep learning models, effectively capturing both global consistency and localized variations within complex 5G network structures.
@@ 1.1.2 Imbalance learning technique
Data-level and algorithm-level methods are the two main categories of class-imbalanced learning techniques. Before building classifiers, data-level approaches preprocess training data to reduce inequality [46]. These strategies include undersampling majority classes and oversampling minority classes. On the other hand, algorithm-level approaches address the issue of class imbalance by modifying the model’s fundamental learning decision-making process. Algorithm-level techniques can be broadly classified into three categories: threshold moving, cost-sensitive learning, and new loss functions [47]. Several approaches have been proposed at the data level. SMOTE [25] is one such technique that generates synthetic minority samples in the feature space by interpolating existing minority samples and their nearest minority neighbors. However, one of the primary limitations of SMOTE is that it creates new synthetic samples without considering the neighborhood samples of the majority classes, which can lead to increased class overlapping and extra noise [48]. Based on the SMOTE principle, many variations have been proposed, such as Borderline-SMOTE [49] and Safe-Level-SMOTE [50], which enhance the original approach by taking majority class neighbors into account. At the same time, Safe-Level-SMOTE creates safe zones to prevent oversampling in overlapping or noisy regions, whereas Borderline-SMOTE only samples the minority samples close to the class borders. To address class-imbalanced challenges, [51] modified the parameters of the model learning process that favored classes with fewer samples. For example, Focal Loss, introduced in [52], allows minority samples to contribute more to the loss function. Additionally, a novel loss function called Mean Squared False Error (MSFE) was proposed in [53] to train deep networks on imbalanced datasets.
@@
This study enables us to handle the class-imbalanced problem on graph data, two new models have been developed recently: We employ some of the settings in GATE-GNN [29] dynamically modifies the weights of various GNN modules and Reweight [54] that focuses on class-incremental learning on dynamic real-world environments on the dual imbalances.
@@
The paper was structured as follows: The notation and preliminaries are introduced in Section (2). In Section (3), we investigate the graph filtering mechanism of the proposed method and the 5G network digital twin datasets. The experiment and evaluation are provided in Section (4). details the datasets and results. We finally concluded in Section (5).
@@ 2 Notations and Preliminaries
We formulated the network failure injection as a classification problem, focusing on learning meaningful embeddings from the generated datasets to increase detection accuracy and identify different failure types. In the network structure, we define a graph $G=(V,E)$ with an adjacent matrix $A=\{e_{vu}|\forall v,u\in V\}\in\mathbb{R}^{N\times N}$, where $N$ represents the number of nodes, letting $G=(V,E)$ represent an input graph with a set of nodes $V$. If $e_{vu}=1$, then $(v,u)\in E$; otherwise, $(v,u)\in E^{\prime}$. We assume the set of feature vectors associated with each network component $v$ are $D$-dimensional represented by $F=\{f_{v}|v\in V\}\in\mathbb{R}^{D}$. We considered the failure points as nodes in the network graph for network failure classification. Each node is associated with various features, such as performance metrics or operational data, reflecting its behavior within the network. The edges connecting pairs of nodes represent the dependencies and relationships between them.
Figure 1: The architectural framework of the proposed technique for imbalanced classification of 5G network of digital twin
The proposed framework in Figure 1 illustrates the proposed model. We applied CF-GNN to examine individual nodes with different types of features (e.g., phys-address, oper-status, etc.). The GNN takes the provided NDT data as input. The GNN retrieves fine-grained information about node interactions and local graph structures, while the Class-Fourier Transform (CFT) efficiently denoizes features while capturing global network attributes. The graph module then processes these inferred relationships to create a multi-layer graph representation. The feature extraction module applies Fourier transform algorithm to convert 5G network data into useful information. To achieve accurate multi-class classification results, the proposed CF-GNN integrate the extracted features with the learned graph structure.
@@
$G=(V,E)$ Graph representing the NDT system
@@
$W_{fc1},b_{fc1}$ Parameters of the GFT linear layer
$W_{mp1},b_{mp1}$ Parameters of the message passing layer

### v2

# Graph Neural Network-based Spectral Filtering Mechanism for Imbalance Classification in Network Digital Twins

@@ Abstract
Graph neural networks are gaining attention in fifth-generation (5G) core network digital twins, which are data-driven complex systems with numerous components. Analyzing these data can be challenging due to rare failure types, leading to imbalanced classification in multiclass settings. Digital twins of 5G networks increasingly employ graph classification as the main method for identifying failure types. However, the skewed distribution of failure occurrences is a significant class-imbalance problem that prevents practical graph data mining. Previous studies have not sufficiently addressed this complex problem. This paper, proposes class-Fourier GNN (CF-GNN) that introduces a class-oriented spectral filtering mechanism to ensure precise classification by estimating a unique spectral filter for each class. This work employs eigenvalue and eigenvector spectral filtering to capture and adapt to variations in minority classes, ensuring accurate class-specific feature discrimination, and adept at graph representation learning for complex local structures among neighbors in an end-to-end setting. The extensive experiments demonstrate that the proposed CF-GNN could help create new techniques for enhancing classifiers and investigate the characteristics of the multiclass imbalanced data in a network digital twin system.
@@
Graph neural network , imbalanced classification , graph representation learning , network digital twin , spectral filtering
@@
[…]
[2]organization= Electronics and Telecommunications Research Institute, postcode=34129, state=Daejeon, country=South Korea
@@ abstract
A novel CF-GNN addresses class-imbalance tasks in multiclass 5G network digital twins datasets.
Eigenvalues and eigenvectors detect the majority and minority classes of the failure types.
Global graph structures and local information affect spectral filtering.
The model performed robustly on the network digital twin-unbalanced datasets.
@@ 1 Introduction
The emergence of fifth-generation (5G) networks and robust core network systems is paramount for the evolution to Industry 5.0. According to recent studies, 5G connections are projected to increase more than 100 times, from approximately 13 million in 2018 to 2.6 billion by 2025 [1]. This growth is expected to continue with sixth-generation networks and the advent of the Industry 5.0 era, characterized by the vision of the Internet of Everything. Classifying network failures for 5G and beyond [2] is essential to optimize network performance. As the backbone of the communication infrastructure, reliable and resilient networks are vital for seamless operation in diverse applications in environments ranging from daily life to industrial settings [3].
The stability, speed, and security of internet connectivity are critical to the reliable functioning of 5G networks [4]. However, disruptions in connectivity can lead to severe consequences. Given the interconnected nature of Internet Service Providers (ISPs), failures originating within a single ISP’s domain can rapidly propagate across the entire network. As such, managing large-scale failures requires significant expertise. To mitigate the impact of such disruptions, rapid and automated anomaly detection is essential. The Border Gateway Protocol (BGP) [5], which interconnects the IP backbone of different ISPs, plays a pivotal role in this context. BGP routers must continuously update internal and external routing information to maintain service continuity. Consequently, the timely identification of hardware and software faults within these routers is crucial for sustaining 5G services [6]. The growing complexity and volume of network traffic further complicate data-driven optimization strategies. In this setting, advanced AI techniques offer promising solutions by shifting the focus from reactive fault handling to proactive fault prediction, enabling preemptive interventions.
Graph classification has emerged as a solution to address the challenges of unbalanced graph classification in applications, including 5G network digital twins [7], fraud detection [8, 9], and, more recently, synthetic oversampling [10], which has gained significant attention. Robust graph neural network (GNN) models must be developed to mitigate biases toward majority classes while ensuring effective generalization for minority classes. Traditional techniques for handling class imbalance, such as oversampling [11] and classifier-based sampling [12], have limitations when applied to graph-structured data because they do not consider the graph structure. Various methods have been proposed to address class imbalance within semi-supervised node classification.
Among the various approaches to address class imbalance in graph learning, synthetic oversampling techniques, including GraphSMOTE [13, 14] and GATE-GNN [15], generate new nodes and establish connections with the existing graph to improve class balance. Furthermore, the Boosting-GNN was introduced [16] to address the classification of imbalanced nodes in graph networks, whereas another study [17] incorporated cost-sensitive learning into the framework. Studies have highlighted that Boosting-GNN improves computational efficiency by assigning higher weights to minority-class nodes during training, particularly in transferring learning applications.
Furthermore, reweighting techniques [18] and resampling methods [19] have been explored to mitigate class imbalance further in GNN-based node classification. In this context, loss functions have also been adapted based on class frequency and node degree, balancing the node distribution while preserving the underlying graph structure. Due to the context of addressing class imbalances in graph learning and focusing on how the loss function behaves in this frequency transformed space becomes critical. To this end, this work addresses the research question: (Q1) How does the consistency of the loss function variance across graphs within the whole graph affect class imbalance, particularly when spectral filtering transforms the graph structure from the time domain to the frequency domain?
Nevertheless, failing to account for the class imbalance when designing GNN models can cause significant performance degradation. If the imbalance is not adequately addressed, the majority class tends to dominate the loss function, causing the trained GNN to overclassify majority class nodes while failing to predict minority class samples accurately. When applied to graphs with skewed class distributions, existing GNN methods frequently overfit the majority class, resulting in suboptimal embeddings for minority class nodes. Addressing this challenge is crucial for improving the generalizability of GNN models and ensuring their effectiveness beyond applications (e.g. 5G network digital twins). However, some studies have attempted to provide an interpretation of GNNs from a spectral perspective [20], suggesting that these methods behave as low-pass filters. More recent work, including GNN based on node adaptive filtering [21, 22], extend this view by considering the influence of node degree, label distributions, and adaptive edge sampling mechanism on filter behaviours. Despite these advances, most GNNs are still primarily designed using a message-passing framework in the spatial domain, with limited exploration into how class-specific information influences spectral filter learning. This raises an important and underexplored question (Q2): How do individual nodes contribute to learning class-specific filter parameters in spectral GNNs? Are certain nodes given more weight during the filter parameter update process, particularly in imbalanced class settings?
This study introduces a novel class-Fourier GNN (CF-GNN) filtering to address imbalance challenges in node-oriented classification and overcome these limitations. This approach employs class-specific spectral filtering, where distinct spectral filters are designed for each class instead of applying a uniform filter to all failure types. This approach ensures that minority-class nodes receive appropriate input without being overshadowed by majority-class nodes, enhancing their representation. The localized filtering mechanism highlights the distinct features of minority class nodes by focusing on specific node characteristics. This targeted approach enables class-specific transformations, ensuring these nodes are adequately represented during training and prediction. In addition, incorporating class-specific spectral awareness in the filtering process reduces the likelihood that minority class nodes are overwhelmed by dominant classes.
Furthermore, to improve classification performance, the proposed model employs class-weighted loss functions, which assign higher penalties to misclassifications of minority class samples. This approach encourages the network to learn more discriminative representations for underrepresented classes. The CF-GNN spectral filtering procedure explicitly integrates a multiclass structure, transforming node features in a class-aware manner using spectral coefficients and eigenvalue-based distinctions between classes. The proposed CF-GNN framework mitigates class-imbalance problems in graph learning, demonstrating substantial improvements over state-of-the-art GNN techniques in the experimental evaluations.
@@
    The proposed CF-GNN employs a spectral filtering mechanism tailored to address class imbalance in multiclass node classification in network digital twins. Using the Fourier transform, CF-GNN captures intricate structural relationships that conventional spatial-based GNNs often neglect. This approach ensures that minority-class nodes are adequately represented, improving classification performance in imbalanced graph scenarios.
@@
    The CF-GNN employs an iterative class-specific spectral filtering approach that dynamically refines eigenvector estimates of the graph structure. This process improves the representation of underrepresented classes by amplifying discriminative characteristics while mitigating the dominance of majority-class nodes in the embedding space, balancing the classification results.
@@
    We introduce the theoretical twin-graph Fourier transform (twin-GFT), a novel multidimensional spectral representation designed to enhance scalability and robustness in failure detection tasks, preserving critical frequency domain characteristics for various types of failures.
@@ 1.1 Related work
Graph learning-driven failure classifications have been gaining attention due to their effectiveness in addressing the massive data volume generated by 5G networks and their ability to understand their complex structural characteristics. This section explores two areas of research: learning-based 5G failure node prediction and imbalanced learning techniques, and the GFT to capture the complex structure of network digital twins.
@@ 1.1.1 Learning-based failure prediction
The primary advantage of applying graph deep learning to network modeling [23] is its data-driven approach, enabling it to capture the complex nature of the real-world networks. Most research [24, 25] has employed fully connected conventional neural networks. However, the primary constraint of these methods is their non-generalizability to alternative network topologies and configurations, such as routing. In this context, more recent studies have proposed advanced neural network models, such as GNNs [26, 27], convolutional neural networks [25], and variational autoencoders [28], for robustness in imbalanced datasets standard in cybersecurity domains. Despite their advancements, these models have different objectives and neglect critical aspects of real-world networks from the twin perspective.
The concept of broadly applying neural networks to graphs has gained significant attention. For example, convolutional networks have been extended to graphs in the spectral domain [29], where filters are applied to frequency modes using GFT. The eigenvector matrix of a graph Laplacian must be multiplied to achieve this transformation. One study [30] parameterized the spectrum filters as Chebyshev polynomials of eigenvalues, resulting in efficient and localized filters and reducing the computational burden. However, a disadvantageous drawback of these spectral formulations is that they are limited to graphs with a single structure because they depend on the fixed spectrum of the graph Laplacian.
In contrast, spatial formulations are not constrained by a specific graph topology. Generalizing neural networks to graphs is studied in two ways: a) given a single graph structure, or labels of individual nodes [31, 29, 32, 33, 30, 34], b) given a set of graphs with different structures and sizes. The goal is to learn predictions of the class labels of the graphs [35, 36, 37]. Inspired by the International Telecommunication Unit Challenge (ITU) ”ITU-ML5G-PS-008: Network Failure Classification Model Using the Digital Twin Network” [38], this research aims to employ the enhanced CF-GNN to analyze interconnected systems and networks in an end-to-end manner. The proposed CF-GNN framework precisely classifies network failure types and accurately identifies failure points by integrating class-oriented spectral filtering. The approach improves accuracy and adaptability compared to conventional deep learning models, effectively capturing both global consistency and localized variations in the network structures.
@@ 1.1.2 Imbalance learning technique
To investigate the underlying causes of failures in digital twin networks, the authors in [39, 5] proposed a fault classification approach based on machine learning (ML). This method involves collecting 41 network-related attributes and employing ML algorithms to analyze these features to identify three primary root causes: CPU overload, node failure, and interface failure. For model evaluation, the study uses three algorithms: Multilayer Perceptron (MLP), Random Forest (RF), and Support Vector Machine (SVM) for failure detection. Figure 1 and Figure 2 represent the class distributions of the two datasets and the imbalance percentage of each failure type. Data-level and algorithm-level methods are the two main categories of class-imbalanced learning techniques. Before building classifiers, data-level approaches preprocess training data to reduce inequality [40]. These strategies include undersampling the majority classes and oversampling the minority classes. In contrast, algorithm-level approaches address the problem of class imbalance by modifying the fundamental learning decision-making process of the model. Algorithm-level techniques can be broadly classified as threshold moving, cost-sensitive learning, and new loss functions [41].
@@
Several approaches have been proposed at the data level. SMOTE [11] generates synthetic minority samples in the feature space by interpolating existing minority samples and their nearest minority neighbors. However, one of the primary limitations of SMOTE is that it creates new synthetic samples without considering the neighborhood samples of the majority classes, which can increase class overlapping and noise [42]. Based on the SMOTE principle, many variations have been proposed (e.g.., Borderline-SMOTE [43] and Safe-Level-SMOTE [44]), enhancing the original approach by considering majority class neighbors. Safe-Level-SMOTE creates safe zones to prevent oversampling in overlapping or noisy regions, whereas Borderline-SMOTE only samples the minority samples close to the class borders. One study [45] modified the parameters of the model learning process that favored classes with fewer samples to address class-imbalance challenges. For example, focal loss [46] allows minority samples to contribute more to the loss function. A novel loss function called mean squared false error [47] was proposed to train deep neural networks on imbalanced datasets. The reviewed literature highlights significant findings from various studies; however, a key research gap remains unexplored and forms the basis of this study. Specifically, existing works have not thoroughly explored the selection of node failure points and optimal features to enhance classifier performance in graph data. Many prior studies concentrate solely on data from physical interfaces, overlooking a broader comparative evaluation of failure detection within network function virtualization (NFV) environments. Addressing this gap, the present work proposes a framework capable of distinguishing failure types from failure points and utilizing the most relevant features.
Figure 1: Class distribution of the failure types in real-world network data (Domain C).
Figure 2: Class distribution of the failure types in network digital twins (Domain A).
1.2 Paper Structure
@@
This study enables us to handle the class-imbalanced graph data problem, and two new models have recently been developed. We employ some of the settings in GATE-GNN [15] to dynamically modify the weights for the GNN modules and the Reweight [48] model, focusing on class-incremental learning in dynamic real-world environments regarding dual imbalances.
@@
The paper is structured as follows: Section (2) introduces the notation and preliminaries. Next Section (3) investigates the graph filtering mechanism of the proposed method on network digital twin datasets. Then, Section (4) presents the experiment and evaluation detailing the datasets and results. Finally, the work is concluded in Section (5).
Figure 3: The components of fifth-generation core network digital twins failure injection settings.
@@ 2 Notations and Preliminaries
We formulate the network failure injection task as a node classification problem, aiming to learn expressive embeddings from rich operational data for accurate detection and differentiation of failure types. The underlying graph representation of the 5G core network is denoted as $G=(V,E)$, where $V$ is the set of network function nodes and $E$ represents the directed or undirected edges that capture operational dependencies and interactions between functions.
The adjacency matrix is defined as $A=\{e_{vu}\mid\forall v,u\in V\}\in\mathbb{R}^{N\times N}$, which encodes the connectivity between nodes, where $e_{vu}=1$ if an edge exists between nodes $v$ and $u$, and $e_{vu}=0$ otherwise. Each node $v\in V$ corresponds to a specific core network function (e.g., AMF, AUSF, UDM) and is associated with a feature vector $f_{v}\in\mathbb{R}^{D}$, capturing operational metrics such as packet counts, memory usage, CPU load, registration events, GTP traffic, and link speeds. Collectively, the feature matrix is denoted as $F=\{f_{v}\mid v\in V\}\in\mathbb{R}^{N\times D}$.
These features are derived from domain-specific categories including network interface statistics (e.g., out-octets, unicast-pkts), performance metrics (e.g., CPU usage, I/O wait), user registration/authentication logs (e.g., RegInitReq, RegInitSucc), and traffic indicators (e.g., GTP.InDataPkt, GTP.OutDataOct). For instance, high out-unicast-pkts values in AMF nodes may indicate control plane congestion, while changes in oper-status in UPF nodes may signal tunnel instability.
Failures are assumed to manifest at the node level, and each node is assigned a class label $y_{v}\in Y$ indicating its operational state (e.g., normal, failed). Misclassifications are associated with node-specific costs based on criticality, with higher penalties for control-plane entities like AMF, AUSF, and UDM due to their impact on system reliability and service continuity.
Figure 4: The architectural framework of the proposed technique for imbalanced classification of the fifth-generation network of a digital twin
Figure 4 illustrates the proposed model framework. We applied CF-GNN to examine individual nodes with various feature types (e.g., phys-address, oper-status, etc.). The GNN takes the provided network digital twin data as input and retrieves fine-grained information about node interactions and local graph structures. The class Fourier transform denoise features while capturing global network attributes. The graph module processes these inferred relationships to create a multilayer graph representation. The feature extraction applies the Fourier transform algorithm to convert the network data into useful information. The proposed CF-GNN integrates the extracted features with the learned graph structure to achieve accurate multiclass classification results.
@@
$G=(V,E)$ Graph representing the network digital twin system